OpenStudy (anonymous):

please find the solution of 4b-3>-1. show work please!! I'm stuck and confused..

OpenStudy (anonymous):

@Mariam.A

OpenStudy (anonymous):

ok 4b-3=-1 first you need to move -3 to the right side to solve for b

OpenStudy (anonymous):

4b=-1+3 4b=2 now you need to eliminate 4 by dividing by 4 b=2/4 b=1/2

OpenStudy (anonymous):

okay. is this finding a solution, or would that be solving it. cause it says find a solution. that's why I'm confused

OpenStudy (anonymous):

it is the same just termed differently so b>1/2 and you can now graph this on the number line

OpenStudy (anonymous):

understand now?

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

and yes!!

OpenStudy (anonymous):

no problem..